OpenStudy (anonymous):

A certain sum is invested for a certain time. It amounts to Rs 400 if the rate is 6% per annum. But when invested at a rate of 2% per annum, it amounts to Rs 250. What is the sum?

OpenStudy (anonymous):

@ganeshie8 quite similar to your name.. :P

OpenStudy (anonymous):

Probably his brother

OpenStudy (anonymous):

400=p*6*t//100 250=p*2*t/100

OpenStudy (anonymous):

P will remain same..

OpenStudy (anonymous):

And t will also be the same..

OpenStudy (anonymous):

To find P, we have to find t also..

OpenStudy (anonymous):

40000=3pt 12500=pt

OpenStudy (anonymous):

Let me think, I am little bit rusty on these topics..

OpenStudy (anonymous):

200/3+125=575/2=

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

i dont know really

OpenStudy (anonymous):

So Am I..

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Not okay,,

OpenStudy (anonymous):

I got something for you..

OpenStudy (anonymous):

how to do this problem

OpenStudy (anonymous):

\[P = \frac{A_1R_1 - A_2R_2}{R_1 - R_2}\]

OpenStudy (anonymous):

when we should use this formula

OpenStudy (anonymous):

Forget this formula.. I will tell you the method..

OpenStudy (anonymous):

Do you know this formula: \[A = (P+ \frac{PRT}{100})\]

OpenStudy (anonymous):

This comes from : A = P + SI \[A = P + \frac{PRT}{100}\] Right??

OpenStudy (anonymous):

dont know

OpenStudy (anonymous):

right

OpenStudy (anonymous):

then

OpenStudy (anonymous):

\[A_1 = P + \frac{6PT}{100}\]

OpenStudy (anonymous):

So: \[400 = P + \frac{6PT}{100} \\250 = P+ \frac{2PT}{100} \\Subtracting \\150 = \frac{4PT}{100} \\PT = 3750\]

OpenStudy (anonymous):

Just check the calculation if I am doing right or not.. You must be having calculator or your genius mind with you.. :P

OpenStudy (anonymous):

Now, can you find SI for any one of the given interest??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[SI_2 = \frac{PRT}{100 }\implies SI_2 = \frac{2 \times 3750}{100} \implies SI_2 = 75\]

OpenStudy (anonymous):

a=p+225

OpenStudy (anonymous):

I found it for second interest ie 2%.. You can also find it for 6%, it is your wish. Amount = P + 75

OpenStudy (anonymous):

It is for second interest rate, when you will do for 1st, you will get the same answer, don't worry about that.. So: \[250 = P + 75 \implies P = 175\]

OpenStudy (anonymous):

175+75=250

OpenStudy (anonymous):

Suppose you find SI for 6%, then you will get : SI = 225.. And A = P + 225: P = 175 again.. It is for 6% case..

OpenStudy (anonymous):

so amount 175 ahh

OpenStudy (anonymous):

can you find time now??

OpenStudy (anonymous):

Here is a shortcut, but you will not follow, just learn the method of doing it, but I am giving it to you.. :)

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

I will write here, and you can do your work, it will take time to write.. :)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

\(\text{If a certaiin sum P in a certain time, amounts to} \; A_1, \text{at the rate of} \; R_1 \%\), \(\text{and the same sum amounts to} \; A_2, \text{at the rate of} \; R_2 \%\), \(\text{then}\) \[P = \frac{A_2 R_1 - A_1R_2}{R_1 - R_2}\] \[T = (\frac{A_2 - A_1}{A_2R_1 - A_1R_2}) \times 100\]

OpenStudy (anonymous):

okay thank you so much

OpenStudy (anonymous):

Only to these type of problems, this formula will apply but I suggest you to learn the method, these formulas are just derived by letting 6% as R_1 and 400 as A_1, so they have become general formulae.. :) Keep it up..